OpenStudy (anonymous):

Solve the system of equations. 5x - 3y = 8 2x + 4y = -2 a.x= -3,y= 1 b.x= 4,y= 4 c.x= 1,y= -1 d.x= 2,y= -3

OpenStudy (anonymous):

have you tried anything?

OpenStudy (anonymous):

i think it is b, but i am not 100% sure

OpenStudy (anonymous):

what makes you think it's b?

OpenStudy (anonymous):

did you work it out?

OpenStudy (anonymous):

what did you do first?

OpenStudy (anonymous):

well the way i was doing it was subtracting, i did it in my head so i really don't think i am correct at all. i never really was taught how to do system of equations, so thats what i need help with the most

OpenStudy (anonymous):

Multiply first equation by 2 and show me what you get..

OpenStudy (anonymous):

10x-6y=16

OpenStudy (anonymous):

Yes.. Now multiply the second equation by 5 and let us see what you got..

OpenStudy (anonymous):

10x+20y=-10

OpenStudy (anonymous):

YES... Now can you subtract both the equations??

OpenStudy (anonymous):

14y=6

OpenStudy (anonymous):

NO.. There is a minus mistake.. 10x - 6y = 16 10x + 20y = -10 On subtracting, 10x and 10x will cancel.. -6y - (+20y) = 16 - (-10) Getting??

OpenStudy (anonymous):

so would it be -26=6? im not understanding this final part

OpenStudy (anonymous):

No.. What you get if you multiply - by - ??

OpenStudy (anonymous):

a positive

OpenStudy (anonymous):

Right.. Then, 16 - (-10) =?? Just mutiply - to -10 and see what you will get..

OpenStudy (anonymous):

so do i multiply -6(20)

OpenStudy (anonymous):

See you are getting confused.. -6y - (20y) = 16 - (-10) This is the final equation.. Now firstly solve for Left Hand Side.. -6y - (20y) = ?? Tell me..

OpenStudy (anonymous):

-26y=26

OpenStudy (anonymous):

Yes now you got it.. Now divide by -26 both the sides and find y..

OpenStudy (anonymous):

1,-1? so c?

OpenStudy (anonymous):

Have faith.. You are right...

OpenStudy (anonymous):

Thank you so much! i appriciate all of your help. i really do

OpenStudy (anonymous):

Welcome dear..